OpenStudy (anonymous):

What is the converse of the following conditional statement? If x = –10, then x2 = 100

OpenStudy (anonymous):

x2would = -20. what?

OpenStudy (the_fizicx99):

I'm assuming that x2 = x^2. -10^2 = 100

OpenStudy (anonymous):

replace the "if" by a "then" and the "then" by an "if"

OpenStudy (anonymous):

If x2 = 100, then x = –10. If x2 mc040-3.jpg 100, then x mc040-4.jpg –10. If x mc040-1.jpg –10, then x2 mc040-2.jpg 100. x2 = 100 if and only if x = –10. these are the answers

OpenStudy (anonymous):

if \(x = –10\) then \(x^2 = 100\) converse if \(x^2=100\) then \(x=-10\)

OpenStudy (anonymous):

first statement it true the converse is false

OpenStudy (anonymous):

thank you